OpenStudy (anonymous):

The sum of the measures of two supplementary angles exceeds the difference of their measures by 116 degrees. Find the measure of each angle

OpenStudy (kamille):

do you know what is the sum of two supplementary angles?

OpenStudy (anonymous):

supplementary angles add up to 180 degrees

OpenStudy (danjs):

Call the angles A and B, A + B = A - B + 116 The sum of the two angles is 116 greater than the difference of the two angles

OpenStudy (danjs):

They are supplementary angles, A + B = 180

OpenStudy (danjs):

You get A + B = 180 180 = A - B + 116 ------------------- A + B = 180 A - B = 64 add the two equations to solve ---------- 2A = 244 ---> A = 122

OpenStudy (danjs):

A = 122 ; B = 64 They are supplementary, and the sum A + B exceeds A - B by 116